OpenStudy (anonymous):

Can someone help me prove sin^2(x)cos^2(x) = (1-cos4x)/theta? I'm having trouble with trigonometry and I can't really grasp the material. Please and thank you

hartnn (hartnn):

\(\cos 2x = 2\cos^2x-1 = 1-2\sin^2x\) so can you find cos^2x =..... ? and sin^2x =.... from this formula ?

OpenStudy (anonymous):

Not really :/

hartnn (hartnn):

\(\cos 2x = 2\cos^2x-1 \) Add 1 to both sides, what u get ?

OpenStudy (anonymous):

1 + cos2x = 2 cos^2 x?

hartnn (hartnn):

thats correct, now divide both sides by 2 what u get ?

OpenStudy (anonymous):

(1+cos2x)/2 = cos^2 x?

hartnn (hartnn):

yes, correct. in the same way, isolate sin^2 x from \(\cos 2x = 1-2\sin^2x\)

OpenStudy (anonymous):

so then cos2x-1 = -2sin^2 x?

hartnn (hartnn):

yes, now divide both sides by -2 what u get ?

OpenStudy (anonymous):

(cos2x-1)/-2 = sin^2 x?

OpenStudy (anonymous):

but I thought you couldn't divide by -2 because it's with the sine?

hartnn (hartnn):

yes, or you can say, (1-cos 2x)/2 = sin^2 x you can divide, there's no problem with that. now substitute both these results in left side of your equation sin^2x . cos^2x =... ?

OpenStudy (anonymous):

what happened to the negative sign that was on the 2?

hartnn (hartnn):

cos2x -1 got turned into 1-cos2x because of that negative sign.

OpenStudy (anonymous):

ohh okay. so ((1-cos^2 x)/2) ((1+cos2x)/2)) ?

hartnn (hartnn):

how cos^2 x ? aren't both of them cos 2x ?

OpenStudy (anonymous):

Darn. Yeah, you're right. I was looking at something else when I was typing

hartnn (hartnn):

so, ((1-cos2 x)/2) ((1+cos2x)/2)) now use \((a-b)(a+b)=a^2+b^2\) so, (1-cos2x)(1+cos2x) =... ?

OpenStudy (anonymous):

1+cos2x-cos2x-2cos^2 x

hartnn (hartnn):

check your last term. its not -2cos^2x

OpenStudy (anonymous):

errr..... -cos^2 x^2? or -cos^2 2x?

hartnn (hartnn):

-cos^2 2x

hartnn (hartnn):

so, you have 1-cos^2 2x according to \(\cos 2x = 2\cos^2x-1\) what will be cos^2 2x=... ?)

OpenStudy (anonymous):

cos2x?

hartnn (hartnn):

\(\cos 2x = 2\cos^2x-1 \\cos 2(2x) = 2\cos^2(2x)-1 \\ \cos^22x =(\cos 4x+1)/2 \) got this ?

OpenStudy (anonymous):

kind of yeah

hartnn (hartnn):

so, whats \(\\~ \\ 1-\cos^22x =1-(\cos 4x+1)/2=..?\)

OpenStudy (anonymous):

cos 2 (2x) = cos ^2 (2x) ?

hartnn (hartnn):

ou had \( ((1-\cos2 x)/2) (1+\cos2x)/2))= (1-\cos^2 2x)/4 \\ =(2-\cos 4x-1)=1-\cos 4x\)

OpenStudy (aravindg):

@treefriendly got it now?

OpenStudy (anonymous):

Sort of, but I'm a bit lost.

OpenStudy (aravindg):

can you tell me where?

OpenStudy (anonymous):

The last part. What happened to the 8 from the initial equation?

OpenStudy (aravindg):

i dont see any 8 here

OpenStudy (anonymous):

Sorry, it's getting late and I've been misreading

OpenStudy (aravindg):

i see ..then i suggest you take rest do it afresh the next day

OpenStudy (anonymous):

I meant what happened to theta that was being divided from 1-cos4x?

OpenStudy (aravindg):

there is no theta here

OpenStudy (anonymous):

prove sin^2(x)cos^2(x) = (1-cos4x)/theta?

hartnn (hartnn):

its not theta, its 8 \(((1-\cos2 x)/2) (1+\cos2x)/2))= (1-\cos^2 2x)/4 \\\\=(1-(\cos4x+1)/2)/4\\ =(2-\cos 4x-1)/8=(1-\cos 4x)/8\)